vmills550:

in triangle FGH, FH = 7 ft., FG = 12 ft., and m

chuckthedonut:

Is there any pictures or answer choices?

mayaluvu:

oh ok

surjithayer:

using cosine formula let Gh=x \[\cos 70=\frac{ 7^2+12^2-x^2 }{ 2\times7\times12 }\] \[168 \cos 70=49+144-x^2\] \[x^2=193-168 \cos 70\] \[x \approx11.64~ft\] using cosine formula \[\cos G=\frac{ 11.64^2+12^2-7^2 }{ 2\times12\times11.64 }\] \[m <G=\cos^{-1} (\frac{ 11.64^2+144-49 }{ 24\times11.64})\] G=?

surjithayer:

G=34.4 degree